OpenStudy (anonymous):

please help me with this problem 12-(1/9)/(4/3)+(1/2)^2 simplify

OpenStudy (anonymous):

Do you know about the order of operations/ PEMDAS?

OpenStudy (anonymous):

yes i can tell you the different answers that i have gotten

OpenStudy (anonymous):

Is this the problem: \[12-\frac{\frac{1}{9}}{\frac{4}{3}}+ (\frac{1}{2})^{2}\] or this: \[\frac{12-\frac{1}{9}}{\frac{4}{3}}+ (\frac{1}{2})^{2}\]

OpenStudy (anonymous):

Looks to me like it'd be the top one.

OpenStudy (anonymous):

@miss texas: I think we'd be better off solving it from scratch here. So tell me this: what is the first thing we're going to do?

OpenStudy (anonymous):

(1/2)^2

OpenStudy (anonymous):

the exponent goes to both terms: \[(\frac{1}{2})^2= \frac{1^2}{2^2}= \frac{1}{4}\]

OpenStudy (anonymous):

Excellent. Nadeem's got that part for us. Is that also what you got?

OpenStudy (anonymous):

oh thats where i am messing up i believe

OpenStudy (anonymous):

Or alternatively, you can do "a half times a half is a fourth" if you know that by heart.

OpenStudy (anonymous):

Well, let's continue working; since there's three of us here, we can catch one another's mistakes. What would you do next?

OpenStudy (anonymous):

We currently have: 12-(1/9)/(4/3)+(1/4)

OpenStudy (anonymous):

give me one second i know that was the first mistake i made reworking the problem

OpenStudy (anonymous):

Sure thing. Take your time.

OpenStudy (anonymous):

flip the second fraction after the division sign?

OpenStudy (anonymous):

Perfect. And then multiply, right?

OpenStudy (anonymous):

yeah but have to get the common denominatior which is 36?

OpenStudy (anonymous):

No need for multiplication. Only need a common denominator for addition and subtraction.

OpenStudy (anonymous):

work on simplifying the fraction: \[\frac{\frac{1}{9}}{\frac{4}{3}}\] The algebra identity is: \[\frac{\frac{a}{b}}{\frac{c}{d}}= \frac{a}{b}\frac{d}{c}= \frac{ad}{bc}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

For example, \[\frac{4}{5}\] times \[\frac{1}{2}\] equals 4/10

OpenStudy (anonymous):

3/36

OpenStudy (anonymous):

so in our case a=1, b=9, c=4, d=3

OpenStudy (anonymous):

Perfect.

OpenStudy (anonymous):

So now we are down to: 12-(3/36)+(1/4)

OpenStudy (anonymous):

But, before we do anything, see any fractions we can simplify?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

none can be simplified so now do i get the common denominator

OpenStudy (anonymous):

Actually, look at 3/36

OpenStudy (anonymous):

Try dividing 36 by 3

OpenStudy (anonymous):

(I'll go ahead and tell you that 36/3 = 12 to save you some work)

OpenStudy (anonymous):

So what does that turn 3/36 into? I know I might seem really pushy, but we've gotta do this to make the next step MUCH easier.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

now you should currently have : \[12-\frac{1}{12}+\frac{1}{4}\]

OpenStudy (anonymous):

Yup, nadine's got it for us.

OpenStudy (anonymous):

Whoops *nadeem

OpenStudy (anonymous):

lol... thats fine

OpenStudy (anonymous):

Alright, so what is 12 in fraction form, misstexas?

OpenStudy (anonymous):

1/12

OpenStudy (anonymous):

Very close. But 1/12 is something else.

OpenStudy (anonymous):

do i get common denominators now?

OpenStudy (anonymous):

No, not yet. Because what you said would imply that two equals 1 one half.

OpenStudy (anonymous):

12/1

OpenStudy (anonymous):

Perfect!

OpenStudy (anonymous):

Alright, *now* we get common denominators. Yay!

OpenStudy (anonymous):

of all three fractions?

OpenStudy (anonymous):

Our denominators: 1, 12 and 4.

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

So, what number do you think we're aiming to turn all three denominators into?

OpenStudy (anonymous):

12 common denominator

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Perfect. Go ahead and tell us what our fractions will look like now.

OpenStudy (anonymous):

146/12

OpenStudy (anonymous):

Excellent! Now, a useful trick: anytime you have a fraction like this: even/even you can divide by 2 to simplify a little.

OpenStudy (anonymous):

and as you can see, your fraction is indeed an even number over an even number. Go ahead and divide both by 2.

OpenStudy (anonymous):

73/6

OpenStudy (anonymous):

Great work! I'll save you time of checking everything and tell you that you've fully simplified, and have the correct answer. Congrats!

OpenStudy (anonymous):

thank you so much for your help are you going to be on much longer

OpenStudy (anonymous):

Sure, I'll see what I can do. Go ahead and post whenever you're ready. And thank YOU for actually staying and replying to our questions. So many folks here just want the answer and don't care about learning. I'm glad you're not one of them.

OpenStudy (anonymous):

I'll probably be heading to bed in 10 minutes. Found any questions for me? If not, just click my profile and "fan" me, and you can check again with me tomorrow or some other day.